Five years ago a mother was twice as old as her son in six years time the sum of their ages will be 82. Find their present ages.
Answer by stanbon(75887)

Five years ago a mother was twice as old as her son.
In six years time the sum of their ages will be 82.
Find their present ages.
--------
Equations:
m-5 = 2(s-5)
m+6 + s+6 = 82
------
Rearrange::
m - 2s = -5
m + s = 70
--------------------
Subtract and solve for "s":
3s = 75
son's age = 25 yrs
-----------
Solve for "m":
m + s = 70
mother's age = 55 yrs
---------
Cheers,
Stan H.